OpenStudy (jjuden):

What is the sine of a 30° angle?

OpenStudy (misty1212):

HI!!

OpenStudy (jjuden):

\[a. \frac{ \sqrt{3} }{ 3 } b. \frac{ \sqrt{3} }{ 2}c. \frac{ 1 }{ 2 }d.\sqrt{3}\]

OpenStudy (jjuden):

hi

OpenStudy (misty1212):

you need a unit circle cheat sheet

OpenStudy (misty1212):

find 30 degrees on the unit circle then find the coordinates of that point the second coordinate is sine

OpenStudy (jjuden):

there are 2 options that will work

OpenStudy (misty1212):

no only one

OpenStudy (misty1212):

sine is the SECOND coordinate

OpenStudy (jjuden):

so 1/2

OpenStudy (misty1212):

yes

OpenStudy (jjuden):

ok thank you :)

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]